Challenge Problem: Momentum and Kinetic Energy An explosion causes an object, initially at rest, to break into two pieces. One of these pieces has twice as much kinetic energy as the other. What is the ratio of their masses? Which one has the larger mass? Is it the one with the larger kinetic energy, or the one with the smaller kinetic energy?
Answer #1

Before explosion the object is at rest, after explosion the two masses 1, T2 move with different velocities v_1,v_2 respectively.

Conserving linear momentum of two masses, 0=m_1v_1+m_2v_2 (or) p1 +p2 = 0  Rightarrow p_1=-p_2 ---(1)

Kinetic energy of mass m1 is K1 = 2m1   P1

Kinetic energy of mass m_2 is 2m2 , using (1) K_2=rac{p_1^2}{2m_2} Rightarrow m_2=rac{p_1^2}{K_2}

The mass m1 has twice as much kinetic energy as the other mass m_2.. That is K_1=2K_2

Hence mass m_2 can be written as m_2=rac{p_1^2}{K_1/2}=2rac{p_1^2}{K_1}=2m_1

That is mass m_2 is double the mass m1.

m1 m2 2

Mass with smaller kinetic energy has larger mass.